24 October 2014 - Mission Day: 6902 - DOY: 297
Recovery
 
 

BLOGS: Recovery Updates after loss of gyros (21 December 1998 - 2 February 1999)

February 2, 1999

    SOHO is back in normal mode!

    The MDI structure programme has already been started and is running. The first EIT picture was obtained at 21:19 UT. It is available at: http://umbra.nascom.nasa.gov/images/EIT_gyroless.gif

    In the meantime, the FPSS off-set has been corrected and the Sun is again centered in the EIT images.
    Check out: http://umbra.nascom.nasa.gov/soho/gyroless.html later tonight (as if Joe has nothing else to do ...).

    Again, my heartiests thanks to everybody who helped in this recovery. What a team!

    Bernhard Fleck

February 1, 1999

    The station keeping maneuver planned for today was successfully executed as planned (delta-V: 32 cm/s). The scheme to balance the torque from the delta-V thruster firings worked well, so no momentum management was necessary after the burn.

    Tomorrow we will return to normal mode with high rate telemetry.

January 30, 1999

    After 40 days, SOHO is out of ESR and im RMW mode, as of 30 January 1999, 17:45 UT! The maneuver went smoothly and almost as planned.

    At a roll rate of about 60 arcsec/s several successful attempts to track a star were made. The effect of the pitch thruster firings to decrease the roll rate, however, was smaller than expected, so the AOCS experts decided to reduce the roll rate by use of the momentum wheels. At a roll rate of 12 arcsec/s the star sensor unit was able to track a star for a sufficient time to allow the spacecraft to be transitioned into RMW mode at 17:45 UT.
    A SSU SEU test was successfully performed.
    A momentum management was successfully performed. The wheel speeds are now 2000/-2000/2000 rpm.
    The present roll attitude is +297.37 deg from ecliptic North. The transition to normal roll attitude has been delayed while the AOCS experts are determining the best method of effecting the roll.

      Transition to FSA and RMW. DSN hot backup, required from step 9b onwards, starts at 16:00 UT. Step Status UT Description 9a. Completed 14:15 Verify configuration for transition from ESR to FSA mode. 9b. Completed 16:02 Trim roll rate to +100 arcsec/sec. 9c. Completed 16:11 Trigger transition from ESR to FSA mode. SOHO in FSA mode 10. Completed 16:17 Initiate SSU map. 11. Completed 16:18 Roll rate monitoring. 12a.Completed 17:44 Wait for valid guide star. 12b.Completed 17:44 Wait for a roll rate lower than 5 arcsec/sec. 12c.Completed 17:45 Trigger transition from FSA to RMW. SOHO in RMW mode 13. Completed 17:49 Observe RMW convergence and SSU locked on star. 14. Completed 18:05 Disable standard monitoring. 15. Completed 19:39 Determine absolute roll attitude: 297.37 deg 16. Completed 18:29 Switch from PROS B to PROS A branch. All activities completed. SSU SEU test in RMW Step Status UT Description 1. Completed 18:41 Verify initial configuration. 2. Completed 18:41 Dump ACU memory @7040. 3. Completed 18:41 Observe ACU patch 5 and 6 parameters. 4. Completed 18:49 Switch to an invalid guide star. 5. Completed 18:49 Monitor results. 6. Completed 18:50 Dump ACU memory @E2A4. 7. Completed 18:50 Dump ACU memory @7040. 8. Completed 18:52 Switch to original guide star. 9. Completed 18:54 Observe ACU patch 5 and 6 parameters. 10. Completed 18:56 Toggle ACU A relay 5. 11. Completed 19:02 Reset SSU SEU flag. All activities completed. Momentum Management in RMW Step Status UT Description 1. Completed 19:14 Verify initial configuration. 2. Completed 19:14 Catalyst bed heaters A on. 3. Completed 19:20 Verify and load macros. Verify/load standard monitoring (ch: 1, 2, 6, 7, 4. Completed 21:05 8, 9, 10, 14, 15, 23, 24, 25, 29, 30, 31, 32, 33, 34, 35). 5. Completed 21:15 LV A OPEN / FCV A ON. 6. Completed 22:12 Command momentum management (2000, -2000, 2000). 7. Finishing at ~23:30 Post maneuver configuration.

        Some acronyms: * ACU: Attitude Control Unit * ESR: Emergency Sun Reacquisition (safe) mode * FCV: Flow Control Valve * FPSS: Fine Pointing Sun Sensor0 * FSA: Fine Sun-pointing Acquisition mode * LV: Latch Valve * PROS: Propulsion Subsystem * RMW: Roll Maneuver Wheels mode * SBM: Stand By Mode * SEU: Single Event Upset * SSU: Star Sensor Unit * WDE: Wheel Drive Electronics

January 29, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    Preparations for the ESR escape maneuver started at 15:00 UT and were successfully completed at 21:25 UT.

      Step Status UT Description ---- ------ ----- ------------ 1. Completed 16:13 Define standard monitoring channels to switch off the ACU and trigger ESR. 2. Completed 17:43 Transition from ACU B to ACU A. 3. Completed 17:51 SSU A switch ON and cool down. 4. Completed 20:20 Uplink ACU patches 5, 6, 7 and wheel spin in SBM patch. 5. Completed 20:52 Uplink SSU patch 2A. 6. Completed 21:11 WDEs 1, 2, 3 switch ON and spin. 7. Completed 21:25 FPSS A switch ON and checkout. 8. As needed N/A Nominal roll rate correction.

    Tomorrow at 14:00 UT we will continue live updates at: http://sohowww.nascom.nasa.gov/operations/maneuver-1999/current.html

      Some acronyms:

      ACU: Attitude Control Unit

      ESR: Emergency Sun Reacquisition (safe) mode

      FCV: Flow Control Valve

      FPSS: Fine Pointing Sun Sensor0

      FSA: Fine Sun-pointing Acquisition mode

      LV: Latch Valve

      PROS: Propulsion Subsystem

      RMW: Roll Maneuver Wheels mode

      SBM: Stand By Mode

      SSU: Star Sensor Unit

      WDE: Wheel Drive Electronics

January 29, 1999 (~15:00UT)

January 28, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    Post burn analysis of the station keeping maneuver performed on 26 January confirmed the good performance indicated by the preliminary data. 4.27 kg of fuel were used. Estimated fuel remaining is 148.4 kg.

    Tomorrow, Friday (29 January) we will start with the preparations for the ESR exit maneuver on Saturday, 30 January, for which 6 hours of hot back up have been requested and have been granted by DSN.

    In order to insure that experts in Europe are available for telephone consultation in the event of an anomaly, as well as giving some of the critical operations folks a well-deserved rest, project management have decided to postpone the momentum and station keeping maneuvers previously planned for Sunday to Monday, 1 February. The new schedule therefore is:

      Friday, 29 January
      Preparation for FSA (ACU B to ACU A, SSU on, patches ACU 5,6,7 & SSU, FPSS A on, wheels on

      Saturday, 30 January
      Transition to FSA and RMW, Roll adjust (SSU SEU test), MMM (increase wheel speeds)

      Monday, 1 February
      MMM (2000/2000/2000), Delta V (cal), Delta V (TBD), MMM

      Tuesday, 2 February
      RMW -> NM transition tests, roll adjust (final), MMM (final)

      Wednesday, 3 February
      Recovery to medium rate telemetry

    Initially we won't follow the Roll Steering Law (RSL) in order to stay on bright stars while we acquire ranging data to obtain a good orbit determination.

    We plan to have "live" status updates on the web on Saturday.

January 26, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    The station keeping maneuver scheduled for today was successfully executed as planned. The target delta-V of 3.8 m/s was achieved. Approximately 4 kg of fuel were used.

January 25, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    The schedule remains as announced on Friday, January 22.

January 22, 1999

    Coding and testing of the ACU software patches to exit ESR without gyro have been successfully completed. In parallel simulations of the ESR exit, momentum management (MoM) and station keeping (SK) in RMW mode were performed, demonstrating their feasibility. The present schedule for next week's activities is:

        Tuesday, 26 January
        Station keeping maneuver (approx. delta-V: 4 m/s)

        D-1
        Spin up wheels and upload ACU patches
        Monitor roll rate

        D
        Exit ESR to RMW (ESR -> FSA -> RMW)

        D+1
        Momentum Management
        Station Keeping Maneuver
        Momentum Mangement
        Roll Trim (TBC)
        Transition to Normal Mode (TBC)

    D day presently is scheduled for Saturday, 30 January. Return to high rate telemetry is planned for the week of February 1 (week 5). Critical support of AOCS experts from ESA/ESTEC and MMS will arrive on Monday, 25 January.

    MoM simulations indicated that it takes about 1 hour per axis, i.e. 3 hours total. Simulations of SK in RMW mode showed that it takes about 2 hours to achieve a delta-V of 1 m/s, and that the wheels saturate after about 6 hours (3 m/s).

    New estimates of the current fuel consumption in ESR with "yaw-braking" is 0.267 kg per day. This is based on assumptions about the thruster duty cycle levels which are estimated at:

      1B/2B duty cycle = 0.06442% (approx. 56 s cumulative firing per day each jet)
      3B/4B duty cycle = 0.05579% (approx. 48 s cumulative firing per day each jet)

      The ratio 0.06443/0.05579=1.15487 yields a net delta-V = 0, i.e. free flight.

January 20, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    The reaction wheels were spun up to test a new method to get an accurate roll rate determination in preparation for the ESR exit. The test was successful and it worked as planned.

January 19, 1999

    The station keeping maneuver scheduled for today was successfully executed as planned in three segments of 20, 30 and 20 min duration. Achieved delta-V was 7.94 m/s for a target of 8.27 m/s. Approximately 9 kg of fuel were used.

    Tomorrow we will spin up the four reaction wheels to get an accurate roll rate determination in preparation for the ESR exit maneuver, still scheduled for next week.

January 15, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    Thanks to a long DSN contact yesterday the daily delta-V increase could be further reduced to 0.10 m/s.

    Yesterday and today a meeting of AOCS experts was held at MMS in Bristol. Good progress was made and a solution to exit ESR was defined. Target date for transition to RMW mode in a success oriented schedule is 28 January.

    The date of the next orbit correction maneuver was confirmed (19 Jan).

January 14, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    The manual yaw damping is continued during DSN contacts.

January 13, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    The manual yaw damping method is continued during DSN contacts. Due to a longer contact yesterday the excess delta-V could be even further reduced from 0.265 m/s per day to 0.15 m/s per day. In order to save hydrazine, we have requested, and have been granted, additional support from DSN. According to the present schedule for the coming days we should get an average of 18.6 hours per day.

January 12, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    The manual yaw damping method is continued. It reduced the orbital disturbance from about 0.45 m/s per day to 0.265 m/s per day.

    The next orbit correction maneuver is planned for next Tuesday, January 19.

January 11, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    The scheme to damp out the yaw oscillation by manually firing single thruster pulses continues to be successfully applied during DSN contacts, which reduces significantly the orbit disturbances caused by the ESR thruster firings.

January 10, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    In preparation to design a fast method to exit the ESR mode, the Star Sensor Unit (SSU) was switched ON, the roll rate manually reduced and a star was tracked for several minutes in both of the two attempts, allowing the determination of the roll rate. The scheme to reduce not only the roll, but also the yaw rate and to increase the pitch rate ("station keeping"), see yesterdays report, is continued.

January 9, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    Ten single thruster pulses were fired manually to keep the yaw rate down (less ESR thruster actuations on thrusters 3 and 4) and to increase the number of ESR thruster actuations on pitch (thrusters 1 and 2). The objectives are to decrease the fuel consumption by the ESR controller and to decrease the excess delta-V as a consequence of the ESR thruster firings.

January 8, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    A test of gyro B confirmed its malfunction.

    A first test to slow down the periodic oscillation in the yaw axis was performed. It was partly successful. The oscillation could be slowed down for about one hour after which it gradually returned to the previous actuation period. During the first hour of the test the yaw thrusters fired 6 pulses compared to the usual 15 pulses. In addition an increased actuation on the pitch control was observed from about 2 pulses per hour to 12 pulses per hour. While the yaw firings push the spacecraft towards the Sun, the pitch firings pull it back, thus helping to reduce the orbit disturbances caused by the ESR firings.

    Approximately 9 kg of hydrazine were used during the orbit correction of yesterday. Approximately 170 kg of fuel are left.

January 7, 1999

    SOHO continues to be stable in ESR with roll control from ground.

    The orbit correction maneuver planned for today was performed successfully in 4 segments of 5, 15, 28 and 15 min burn time, respectively. The first 3 segments were cold by about 20% which necessitated a fourth segment.

January 6, 1999

    The ESR triggered on 1998 December 21 at 17:49 UT continues. In this mode, the ESR hardware autonomously fires the pitch and yaw thrusters to maintain Sun pointing (and, thus, solar panel illumination). Roll control is "open loop:" roll thruster firings are commanded from the ground once or twice per day to keep the roll rate below 0.3 0/s.

    Deep Space Network hot backup support has been confirmed for the stationkeeping maneuver (the first ever attempted while in ESR) scheduled for today, 1999 January 7 starting at 18:10 UT.

    This maneuver is designed to compensate for the orbital velocity changes created by the frequent yaw thruster firings. The maneuver burns will be conducted in three parts: a short burn (~ 5 minutes) followed by a check of downlink Doppler data, then burns of roughly 15 and 28 minutes each, separated by 20 - 30 minutes to check roll rate crosstalk from the burns (again using Doppler data). Any roll rate above 0.3 0/s will be compensated for by ground commanded thruster firings.

    We would like to thank our colleagues on the Ulysses mission, who have sacrificed some of their DSN support to allow hot backup for the maneuver.

December 22, 1998

    SOHO went into an ESR (Emergency Sun Reacquisition) on 21 December 1998, 17:49 UT, during the setup phase of a planned momentum management and station keeping maneuver, just as the spacecraft was meant to enter a gyro-dependent attitude mode (RMW).

    Indications are that the third and last gyro has been lost. SOHO is now in an ESR "safe hold" mode without the use of gyros to measure deviations from its proper orientation. Sun-pointing attitude is maintained by intermittent thruster firings which are actuated and controlled by on board sun sensors. The roll rate of the spacecraft is controlled using open-loop (ground-commanded) thruster firings. SOHO will remain in this mode until new software for a gyroless mode of operation has been developed.

    Due to insufficient roll control, SOHO cannot use its high gain antenna. The low gain omni-directional antenna is being used instead. Therefore no science data can be transmitted.

    Since the thrust from the ESR firings pushes the spacecraft towards the Sun, periodic thruster firings to maintain the proper orbit are also required. This is called station keeping. Preliminary estimates indicate that SOHO is consuming its fuel at a rate of about 7 kilograms per week as a result of these activities. As of end of December 1998, SOHO had about 180 kg of fuel left.

    The SOHO team requested, and has been granted, additional support from the Deep Space Network to support these activities.

December 21, 1998

    SOHO is in safe mode. The spacecraft entered safe mode today at 17:49 UT. Preliminary indications are that the remaining gyro may have failed.

 
 

• European Site • US Site

 

Feedback & Comments: SOHO Webmaster

Last modification:

SOHO
and
SOHO
SOHO is a project of international cooperation between